### v1

# Collective dynamics versus entanglement in quantum battery performance

Abstract
Identifying the physical origin of enhanced charging performance in many-body quantum batteries is a key challenge in quantum thermodynamics. We investigate whether improvements in stored energy and instantaneous charging power arise from genuine quantum correlations or from coherent collective dynamics that are not intrinsically quantum. We compare the time evolution of energetic quantities with a hierarchy of information-theoretic measures probing bipartite, tripartite, and further-partite correlations. Across different battery–charger configurations, we find a consistent temporal ordering in which the instantaneous power peaks before the buildup of strong quantum correlations, indicating that peak charging is dominated by coherent transport, while entanglement and scrambling develop at later times. Furthermore, charging protocols based on $\kappa$-local interactions are examined under both unconstrained and norm-constrained (fair) settings, enabling a clear distinction between classical scaling effects and genuine collective enhancements. Increasing the interaction order or the participation number does not automatically translate into higher charging power. Instead, the performance is primarily dictated by how many particles actually become mutually correlated and contribute to entanglement. Fully collective interactions provide a genuine advantage because all particles participate coherently, whereas partially extended interaction schemes fail to monotonically increase the number of effectively interacting particles, and therefore do not guarantee improved charging efficiency.
I Introduction
Quantum batteries (QBs) are devices engineered to store and deliver energy through controlled quantum dynamics and have emerged as a central platform in quantum thermodynamics and quantum technologies. Importantly, QBs can be realized using quantum many-body systems, such as interacting spin chains, where collective effects, many-body interactions, and correlations can enhance charging power and lead to nontrivial scaling behaviors [alicki2013entanglement, binder2015quantacell, campaioli2017enhancing, ferraro2018high, friis2018precision, andolina2019quantum, andolina2019extractable, campaioli2018quantum, le2018spin, lyon2020many, kurizki2020quantum, rodriguez2021collective, zhang2022longrange, giuliani2022collective, zanardi2022scrambling]. Beyond conventional spin-chain architectures, QBs have been investigated in topological systems as well as in various spin-based models subject to different types of external driving or interactions, including Floquet-engineered setups, kicked-spin models, and long-range interacting frameworks. These studies highlight the roles of periodic driving, operator spreading, and topological protection in the storage and transfer of energy [lu2025topological, Puri2024Floquet, Romero2025].
Moreover, QBs have been extensively investigated in open quantum system scenarios, where environmental effects such as dissipation, decoherence, and engineered reservoirs strongly influence charging efficiency, stability, and extractable work. These studies reveal phenomena including noise-assisted and dissipative charging, reservoir-mediated energy transfer, non-Markovian memory effects, thermal and composite environment influences, and fundamental power bounds [Ferraro2018, Andolina2019, GarciaPintos2020, Sharma2022, Quach2020, Tabesh2020EnvMediated, Morrone2022NonMarkov, EnhanceOpenQB2023, Yao2025Reservoir, QuantumErgotropy2025, NonMarkovNSpin2025, Xu2023CompositeEnv, Zhang2024OpenInteractive, Zhang2024OpenEnvPerformance, EntropicUncertaintyQB2022]. Collectively, these works emphasize that both the internal structure of the quantum battery and the surrounding environment critically determine charging performance.
From a structural perspective, quantum battery architectures are commonly categorized into two paradigms: (i) interacting batteries charged by noninteracting systems [le2018spin, ghosh2020enhancement, zhao2021quantum], and (ii) noninteracting batteries charged by interacting many-body chargers [rossini2020quantum, ferraro2018high, julia2020bounds]. In both cases, charging performance depends sensitively on the Hamiltonian structure, interaction range, and normalization, with rigorous bounds linking charging power to the Hamiltonian norm [julia2020bounds, abdelkhalek2019].
A central goal in quantum battery research is to identify genuinely quantum mechanisms such as coherence, entanglement, and collective dynamics that can enhance stored energy, extractable work, and especially charging power beyond classical limits [baumgratz2014quantifying, horodecki2009quantum]. Early studies demonstrated that global collective driving can enhance instantaneous charging power relative to parallel local protocols [alicki2013entanglement, binder2015quantacell, campaioli2017enhancing], motivating the idea that entanglement and multipartite correlations might serve as key resources for fast charging. However, subsequent investigations clarified that entanglement is neither necessary nor sufficient for charging enhancement [andolina2019quantum, andolina2019extractable, julia2020bounds, kamin2020entanglement, shi2022entanglement], as maximal charging power is often reached prior to the development of strong multipartite correlations [kamin2020entanglement, shi2022entanglement, perarnau2015extractable].
Recent attention has focused on the role of dynamical complexity, including quantum chaos, operator spreading, and scrambling. Studies based on Floquet engineering, kicked-Ising models, and long-range interactions indicate that rapid operator growth can significantly boost charging performance [Puri2024Floquet, Romero2025]. Likewise, works on information scrambling show that fast propagation of quantum information, characteristic of nonintegrable systems, can optimize energy transfer in many-body setups [nahum2018operator, vonkeyserlingk2018operator, liu2018scrambling, swingle2018unscrambling], suggesting that dynamical properties, rather than static entanglement alone, are decisive in determining charging power.
A key challenge in the study of many-body quantum batteries is identifying the physical origin of enhanced charging performance. In particular, it is crucial to understand whether improvements in stored energy and instantaneous charging power arise from genuinely quantum correlations or from coherent collective dynamics that are not intrinsically quantum. To address this question, we analyze the temporal evolution of energetic quantities alongside a hierarchy of information-theoretic measures probing bipartite, tripartite, and multipartite correlations. Across different battery–charger configurations, we find a consistent temporal ordering: the instantaneous power peaks before the buildup of strong quantum correlations, indicating that coherent transport dominates peak charging, while entanglement and scrambling develop at later times.
We further investigate charging protocols based on $\kappa$-local interactions under both unconstrained and norm-constrained (fair) conditions, which allows us to separate classical scaling effects from genuine collective behavior. Our analysis shows that increasing the interaction order or participation number alone does not guarantee enhanced charging power. Instead, the performance of the battery is governed by the coherence and structure of the interaction network: fully collective interactions provide a genuine advantage, whereas partially extended interaction schemes can suppress power due to competing correlation pathways. These findings clarify the distinct roles of coherence, entanglement, and interaction structure in optimizing the charging performance of many-body quantum batteries.
The paper is organized as follows. Section II introduces the quantum battery and charger Hamiltonians and defines measures for bipartite, tripartite, and multipartite entanglement. Section III presents the charging dynamics for both configurations—interacting battery with noninteracting charger, and noninteracting battery with interacting charger—and examines the effects of $\kappa$-local interactions and interaction range on energy transfer and correlations. Section IV summarizes the main findings and concludes the study.
II set up
Quantum Battery and Charger Hamiltonian
We consider a quantum battery consisting of NN noninteracting spin-$\frac{1}{2}$ particles. The corresponding battery Hamiltonian is defined as
$\hat{H}_{B}=h_{z}\sum_{j=1}^{N}\hat{\sigma}_{j}^{z},$ (1)
[…]
To initiate the charging process, we switch on the interactions among the battery spins at time $t=0$. Additionally, we include an external field applied in the $x$-direction, which further drives the charging dynamics. The resulting evolution is governed by the interacting spin Hamiltonian
$\hat{H}_{C}=\sum_{\alpha=x,y,z}\left(J_{\alpha}\sum_{j=1}^{N-\kappa+1}\left[\prod_{\mathrm{k}}^{\kappa}\hat{\sigma}_{j+\mathrm{k}-1}^{\alpha}\right]\right)+h_{x}\sum_{j=1}^{N}\hat{\sigma}_{j}^{x},$ (2)
where $J_{\alpha}$ are the nearest-neighbor interaction strengths in the $\alpha$-direction, and $h_{x}$ is the magnetic field applied to each spin in the x-direction. $\kappa$ quantifies the number of interacting spins, representing the degree of $\kappa$-locality. By varying $\kappa$, we examine how higher-order spin correlations influence charging performance and the generation of multipartite entanglement. This Hamiltonian simultaneously induces spin correlations and drives the energy injection that charges the battery. The charging dynamics of the battery are governed by the Hamiltonian
$\hat{H}=\hat{H}_{B}+\hat{H}_{C},$ (3)
Energy Storage and Power Dynamics
To characterize the performance of a quantum battery, we focus on key quantities related to energy storage and its rate of transfer. These quantities allow us to analyze how efficiently the battery is charged over time and the role of quantum correlations in this process.
Stored Energy:
The average energy stored in the battery at a given time $t$ is defined as
$W(t)\equiv\Delta E(t)=\mathrm{Tr}\big[\hat{H}_{B}\hat{\rho}(t)\big]-\mathrm{Tr}\big[\hat{H}_{B}\hat{\rho}(0)\big],$ (4)
where $\hat{\rho}(t)=\hat{U}(t)\hat{\rho}(0)\hat{U}^{\dagger}(t)$ is the time-evolved state of the battery under the Hamiltonian $\hat{H}$. Here, $\hat{\rho}(0)$ is the initial state of the battery, typically chosen as the ground state of $\hat{H}_{B}$, and $\hat{U}(t)=e^{-i\hat{H}t}$ is the unitary evolution operator. The quantity $W(t)$ measures the net increase in the battery’s energy due to the charging process, providing a direct quantification of energy storage.
Instantaneous Power:
The instantaneous power captures the rate at which energy is being transferred to the battery at a specific moment in time. It is defined as the time derivative of the stored energy:
$P_{i}(t)=\frac{dW(t)}{dt}.$ (5)
In analyzing the dynamics of quantum batteries, the instantaneous power provides a detailed, time-resolved view of energy transfer, in contrast to the average power, which captures only the overall energy change over a finite interval. Focusing on $P_{i}(t)$ allows us to identify the moments of maximal energy injection and to examine their correlation with the growth of quantum correlations in the battery. By comparing $P_{i}(t)$ with bipartite (concurrence and bipartite entanglement entropy), tripartite (tripartite mutual information), and multipartite entanglement measures (quantum Fisher information and average bipartite entanglement entropy), we can assess whether periods of rapid charging coincide with the buildup of entanglement. This approach establishes a physically meaningful connection between the temporal structure of energy flow and the emergence of genuine quantum effects, distinguishing enhancements arising from coherent many-body interactions from those due purely to classical energy scaling. In the following, we introduce and define all the entanglement measures that are utilized in this study.
Entanglement Measures
Concurrence
Concurrence is one of the most widely used measures of bipartite entanglement for two-qubit systems [wootters1998entanglement]. It provides a quantitative value ranging from 0 for separable (uncorrelated) states to $1$ for maximally entangled states. In the context of quantum batteries, concurrence is useful for quantifying the entanglement between specific pairs of spins, for example, the first and last spins of a chain, which can reveal the formation of long-range correlations during the charging process.
For a two-qubit density matrix $\rho$, we define the spin-flipped matrix as $\tilde{\rho}=(\sigma_{y}\otimes\sigma_{y})\rho^{*}(\sigma_{y}\otimes\sigma_{y})$, where $\rho^{*}$ is the complex conjugate in the computational basis and $\sigma_{y}$ is the Pauli-$y$ matrix. Using this, we construct $R=\rho\tilde{\rho}$. Let $\{\mu_{i}\}_{i=1}^{4}$ be the eigenvalues of $R$ (which are always real and nonnegative), and let their square roots be arranged in decreasing order: $\lambda_{1}\geq\lambda_{2}\geq\lambda_{3}\geq\lambda_{4}\geq 0$. The concurrence is then given by
$C(\rho)=\max\Big(0,\lambda_{1}-\lambda_{2}-\lambda_{3}-\lambda_{4}\Big).$
Concurrence provides insight into how entanglement between two specific subsystems evolves in time. In quantum batteries, it allows us to track the onset of correlations that can facilitate energy transfer, and to determine whether strong bipartite entanglement is required for achieving high charging efficiency.
Bipartite Entanglement Entropy
Bipartite entanglement entropy (BEE) quantifies the degree of entanglement between a subsystem $X$ and its complement $X^{c}$ in a many-body system [Calabrese2005, Page1993]. It is computed from the entropy of the reduced density matrix associated with the subsystem,
$S_{X}=-\mathrm{Tr}_{X}\left[\hat{\rho}_{X}\log\hat{\rho}_{X}\right],$ (6)
where $\hat{\rho}_{X}=\mathrm{Tr}_{X^{c}}\left[\hat{\rho}\right$] is obtained by tracing out all degrees of freedom outside $X$. Here, $\hat{\rho}$ denotes the density matrix of the full system, which we consider to be a pure state, $\hat{\rho}=|\psi_{0}\rangle\langle\psi_{0}|$, with $|\psi_{0}\rangle$ representing the ground state of the battery Hamiltonian $\hat{H}_{B}$.
The BEE captures the total amount of quantum correlations shared between a chosen subsystem and the rest of the system. In the context of quantum batteries, it reveals how local parts of the system become entangled with the rest during the charging process, which can affect how efficiently energy is stored and distributed across the battery.
Tripartite Mutual Information
To study correlations among three subsystems simultaneously, we use the tripartite mutual information (TMI), defined as [iyoda2018scrambling, seshadri2018tripartite, aleiner2016microscopic, xu2022operator, zhang2020information, hosur2016chaos, von2018scrambling, Shukla2025ScramblingIsing]
$I_{3}(X:Y:Z)=I_{2}(X:Y)+I_{2}(X:Z)-I_{2}(X:YZ),$ (7)
where $I_{2}(X:Y)=S_{X}+S_{Y}-S_{XY}$ is the bipartite mutual information. Here, $S_{X}$ is the von Neumann entropy of subsystem $X$, and $S_{XY}$ corresponds to the joint entropy of subsystems $X$ and $Y$.
TMI captures the interplay of correlations among three subsystems, identifying whether correlations are shared redundantly, synergistically, or exclusively between subsets. In quantum batteries, TMI helps us understand how energy and information propagate through multiple subsystems, revealing the emergence of complex entanglement patterns that go beyond pairwise correlations.
Quantum Fisher Information
[…]
$F_{Q}[\hat{H}_{B}]=4\,(\Delta\hat{H}_{B})^{2}=4\Big(\langle\psi_{0}|\hat{H}_{B}^{2}|\psi_{0}\rangle-\langle\psi_{0}|\hat{H}_{B}|\psi_{0}\rangle^{2}\Big),$ (8)
[…]
The QFI quantifies how sensitive the battery state is to unitary transformations generated by $\hat{H}_{B}$, and larger QFI indicates stronger collective correlations across spins, which can enhance energy storage and transfer in quantum batteries.
Furthermore, QFI can witness multipartite entanglement. For an $N$-spin battery, if the inequality
$F_{Q}[\hat{H}_{B}]\leq\Big\lfloor\frac{N}{\kappa}\Big\rfloor\kappa^{2}+\Big(N-\Big\lfloor\frac{N}{\kappa}\Big\rfloor\kappa\Big)^{2}$ (9)
[…]
Average Bipartite Entanglement Entropy
While BEE measures correlations for a single partition, the average bipartite entanglement entropy (ABEE) provides a global view of entanglement across the system [Polkovnikov2011, Riera2012]. It is computed by averaging the BEE over all contiguous bipartitions:
[…]
Figure 1: Noninteracting battery charged by an interacting Hamiltonian. Bipartite entanglement $(a_{1},b_{1},c_{1})$: Time evolution of the instantaneous power $P_{i}$, concurrence between the first and last spins $C(1,N)$, and bipartite entanglement entropy (BEE) $S_{N/2}$. Tripartite entanglement $(a_{2},b_{2},c_{2})$: Time evolution of $P_{i}$ and tripartite mutual information (TMI) $I_{3}$. Multipartite entanglement $(a_{3},b_{3},c_{3})$: Time evolution of $P_{i}$, quantum Fisher information (QFI) $F_{Q}$, and average entanglement entropy (AEE) $\overline{S}_{N/2}$. Panels correspond to: $(a_{1},a_{2},a_{3})$ $J_{x}=1$, $J_{y}=J_{z}=0$; $(b_{1},b_{2},b_{3})$ $J_{x}=J_{y}=1$, $J_{z}=0$; $(c_{1},c_{2},c_{3})$ $J_{x}=J_{y}=J_{z}=1$. Number of spins: $N=8$ and $h_{x}=h_{z}=1$.
III Results
The central aim of our study is to establish whether the enhancement of stored energy in a many-body quantum battery is driven by the buildup of quantum correlations. To address this question we compare the time evolution of energetic quantities (stored energy and instantaneous power) with a hierarchy of information-theoretic measures that probe correlations at different scales: entanglement of bipartition measure (concurrence and BEE), entanglement of tripartition measure (TMI), and entanglement of multipartite measures (QFI and the ABEE). By juxtaposing these temporal profiles we identify how energy transfer, correlation generation, and scrambling are interrelated in the charging process and extract the underlying physical mechanisms.
Charging Rate and Entanglement with $2$-local Interactions
Noninteracting Battery and Interacting Charger
We first consider a configuration in which the battery Hamiltonian, $H_{B}$, represents a noninteracting system, and the same physical system also serves as the charger through the Hamiltonian $H_{C}$. The interaction terms are switched on only during the charging interval $0,\tau]$, at which point the charger becomes active and drives the dynamics. Its internal couplings mediate the energy flow into the battery and simultaneously generate quantum correlations. In this framework, the noninteracting battery acts as a passive energy receptor, allowing us to clearly isolate and analyze the role of charger-induced correlations in the charging process.
Figure 2: Interacting battery charged by an noninteracting Hamiltonian. Bipartite entanglement $(a_{1},b_{1},c_{1})$: Time evolution of the instantaneous power $P_{i}$, concurrence between the first and last spins $C(1,N)$, and bipartite entanglement entropy (BEE) $S_{N/2}$. Tripartite entanglement $(a_{2},b_{2},c_{2})$: Time evolution of $P_{i}$ and tripartite mutual information (TMI) $I_{3}$. Multipartite entanglement $(a_{3},b_{3},c_{3})$: Time evolution of $P_{i}$, quantum Fisher information (QFI) $F_{Q}$, and average entanglement entropy (AEE) $\overline{S}_{N/2}$. Panels correspond to: $(a_{1},a_{2},a_{3})$ $J_{x}=1$, $J_{y}=J_{z}=0$; $(b_{1},b_{2},b_{3})$ $J_{x}=J_{y}=1$, $J_{z}=0$; $(c_{1},c_{2},c_{3})$ $J_{x}=J_{y}=J_{z}=1$. Number of spins: $N=8$ and $h_{x}=h_{z}=1$.
We investigate the relationship between the battery charging rate and bipartite correlations. To quantify bipartite correlations, we evaluate two complementary measures: the concurrence and the BEE. These quantities are computed for three representative charger interaction configurations: ($a_{1}$) $J_{x}=1$, $J_{y}=0$, $J_{z}=0$; ($b_{1}$) $J_{x}=1$, $J_{y}=1$, $J_{z}=0$; and ($c_{1}$) $J_{x}=1$, $J_{y}=1$, $J_{z}=1$. In all cases, the field strength is fixed to $h_{x}=1$. Concurrence quantifies pairwise entanglement between two qubits; here, we focus on the first and last spins to probe long-range correlations. In contrast, the BEE measures the entanglement between a chosen subsystem and its complement through the BEE of the reduced density matrix. In our calculations, we consider equal bipartitions of the spin chain. For a meaningful comparison, all three quantities are normalized by their respective maximum values, allowing us to identify the time regimes in which each quantity attains its peak.
During the initial charging stage the instantaneous power $P_{i}(t)$ rises rapidly as the charger injects excitations into the battery; this rise is driven by coherent excitation transfer enabled by the charger’s internal couplings. Concurrence and BEE also increase during this phase, indicating that entanglement participates in the energy transfer process. Crucially, however, the peak of instantaneous power is reached before concurrence and BEE attain their maxima (see Fig. 1 $(a_{1},b_{1},c_{1})$). Concurrence typically grows faster than BEE and peaks earlier than BEE, but both lag the power peak. Physically, this temporal ordering shows that short-time coherent transport, which governs peak power, precedes the formation of fully developed bipartite entanglement across the system. Entanglement is therefore a necessary resource for enhanced charging, but it is not the sole determinant of the timing of maximal power: the charger’s coupling strengths and the coherent transfer channels set the time scale for the power peak, whereas entanglement continues to build up afterwards as correlations spread.
To probe correlations beyond pairwise contributions, we evaluate the TMI, which quantifies how information and correlations are distributed among three subsystems. For this calculation, the spin chain of length $N$ is partitioned into four segments of size $N/4$ in total, and the initial state is taken to be the ground state of the battery Hamiltonian. This ensures consistency between the injected energy and the resulting correlation production during the charging process. The TMI increases substantially only after the early-time rise of energy and power, and its maximum occurs at later times compared to energetic quantities [Fig. 1 $(a_{2},b_{2},c_{2})$)]. This delayed growth of TMI reveals that three-way and higher-order correlations require additional time to form: pairwise coherences must first be established and propagated before genuine three-body information sharing emerges. From a physical standpoint, this hierarchical buildup reflects the propagation and interference of excitations through the interacting charger, leading ultimately to delocalized correlation structures (scrambling) that are slower to develop than local energy transfer.
To characterize genuinely global correlations we evaluate the QFI and the ABEE. The QFI serves as a witness of multipartite entanglement and quantifies the state’s collective sensitivity to global parameter shifts; ABEE captures the typical entanglement across all contiguous bipartitions of the chain. QFI and ABEE grow more slowly than the instantaneous power and reach their maxima at later times [Fig. 1 $(a_{3},b_{3},c_{3})$)]. This systematic delay indicates that while multipartite entanglement is ultimately established as energy and local correlations spread, the regime of maximal charging power is dominated by coherent, low-order collective modes rather than by fully developed global entanglement. High late-time ABEE values signify that entanglement becomes extensively distributed across the system only after most of the energy transfer has already occurred; hence, multipartite entanglement functions more as a stabilizing resource for stored energy at long times than as the immediate driver of peak power.
Energy injection from an interacting charger initially excites coherent transport channels that rapidly transfer energy into the battery and produce a pronounced peak in instantaneous power. The formation of quantum correlations pairwise, tripartite, and finally multipartite follows this energetic surge because correlations require the accumulation of relative phases and interaction-mediated spreading of excitations. Consequently, entanglement and scrambling develop on longer timescales than energy transfer. Therefore, although entanglement is essential to realize quantum advantages in charging (for instance by enabling nonclassical interference pathways or collective enhancement), there is no one-to-one temporal correspondence between entanglement maxima and the peak of instantaneous power. Instead, optimal charging occurs in an intermediate regime where coherence and modest entanglement combine to maximize power, while more extensive entanglement emerges later and contributes to the global reorganization and stabilization of the stored energy.
Interacting Battery and Noninteracting Charger
We now consider the reverse configuration, where the battery is interacting and the charger is noninteracting, effectively interchanging the roles of the two subsystems studied earlier. This setup allows us to assess how internal interactions within the battery itself influence the charging process and the buildup of correlations. Similar to the previous case, we monitor the time evolution of the instantaneous power, stored energy, and various entanglement measures to understand how the interacting battery dynamics unfold.
For bipartite correlations, we again compute the concurrence and the BEE. Across all interaction configurations, the concurrence exhibits a qualitatively different temporal behavior compared to the previous configuration: it starts with a finite value at $t=0$ and subsequently decreases as the instantaneous power increases [Fig. 2 ($a_{1},b_{1},c_{1}$)]. For the cases $J_{x}=1$, $J_{y}=J_{z}=0$ and $J_{x}=J_{y}=1,J_{z}=0$, the BEE grows more gradually than the instantaneous power and reaches its maximum at later times, consistent with the general trend observed in the interacting-charger setup [Fig. 2 ($a_{1},b_{1}$)]. However, when all interaction terms are present ($J_{x}=J_{y}=J_{z}=1$), the BEE remains nearly constant throughout the evolution, suggesting a suppression of entanglement dynamics despite strong coupling among the spins [Fig. 2 ($c_{1}$)]. Notably, the BEE maximum, when present, still appears after the peak in instantaneous power [Fig. 2 $(a_{1},b_{1},c_{1})$], maintaining the same temporal hierarchy as in the previous case. The early-time surge in power thus originates from rapid coherent excitation transfer between the noninteracting charger and the interacting battery, while the battery’s internal interactions primarily act to redistribute correlations locally. This results in a delayed and modest enhancement of pairwise entanglement without significantly altering the overall bipartite correlation structure of the system. However, an important qualitative difference emerges: while the instantaneous power increases, the quantum coherence in the battery decreases. This indicates that the interacting battery tends to convert quantum coherence into energetically useful, more classical population distributions during the charging process [Fig. 2 $(a_{1},b_{1},c_{1})$].
We next examine higher-order correlations using the TMI. For the cases $J_{x}=1$, $J_{y}=J_{z}=0$ and $J_{x}=J_{y}=1,J_{z}=0$, TMI exhibits a slower growth rate compared to energetic quantities, reaching its maximum at later times [Fig. 2 $(a_{2},b_{2})$]. This delay indicates that the emergence of genuine multipartite correlations is not instantaneous, even when the battery possesses intrinsic interactions. Instead, these correlations develop progressively as excitations propagate and interfere across different regions of the interacting battery. At early times, the charging process is dominated by localized excitations and short-range coherence, while nonlocal correlation sharing becomes prominent only after the system approaches energetic saturation. However, when all interaction terms are present ($J_{x}=J_{y}=J_{z}=1$), similar to the BEE, the TMI also remains constant throughout the evolution [Fig. 2($c_{1}$)]. This behavior can be attributed to the high degree of symmetry in the fully isotropic spin interaction. The isotropic Heisenberg dynamics preserves the global correlation structure of the system, restricting the redistribution of information among different subsystems. As a result, multipartite correlations do not grow or decay during the evolution, leading to a time-invariant TMI despite the ongoing unitary dynamics.
[…]
We observe a strikingly similar behaviour of the instantaneous power in both configurations: a noninteracting battery charged with an interacting charger, and an interacting battery charged with a noninteracting charger (compare Fig. 1 and Fig. 2). The instantaneous power of the quantum battery is defined as $P_{i}(t)=\frac{d}{dt}\langle\hat{H}_{B}\rangle=i\langle[\hat{H},\hat{H}_{B}]\rangle$, where the total Hamiltonian is taken as $\hat{H}=\hat{H}_{B}+\hat{H}_{C}$. Since $\hat{H}_{B},\hat{H}_{B}]=0$, the power is governed entirely by the commutator $P_{i}(t)=i\langle[\hat{H}_{C},\hat{H}_{B}]\rangle$. Thus, the instantaneous power originates purely from the noncommutativity between the battery Hamiltonian and the charger Hamiltonian. In the case of a noninteracting battery with an interacting charger, the interacting nature of $\hat{H}_{C}$ ensures that $\hat{H}_{C},\hat{H}_{B}]\neq 0$, leading to a nontrivial power profile. Conversely, when the battery is interacting and the charger is noninteracting, the structure of $\hat{H}_{B}$ now plays the dominant role in producing a similar nonvanishing commutator. In both situations, the source of power is effectively the same mathematical object, the commutator between the two subsystems, resulting in qualitatively similar early-time behaviour of the instantaneous power. The differences between the two scenarios are expected to manifest only at later times, where the interacting Hamiltonian induces many-body correlations and enhanced entanglement dynamics in the system.
Figure 3: Unfair charging. (a) Parallel charging: Time evolution of the stored energy $\Delta E$ for a noninteracting battery, where each spin is charged independently by a local transverse field of strength $h_{x}=1$ and $h_{x}=N$ (with all spin-spin interactions set to zero). (b) Global charging: Time evolution of the stored energy $\Delta E$ for the same noninteracting battery, charged by an all-to-all interacting spin charger with interaction strengths $J_{x}=1$ and $J_{x}=N$, in the absence of any external transverse field. $N=8$ (open boundary conditions), with $h_{z}=1$ fixing the battery energy scale.
Charging Dynamics under $\kappa$-Local Interactions: Role of Interaction Structure
Understanding the role of interaction structure in quantum battery charging is central to identifying genuine quantum advantages. Early theoretical studies have demonstrated that collective interactions can enhance charging power compared to parallel protocols [campaioli2017enhancing]. However, it was later clarified that such advantages may arise trivially from increasing the overall energy scale of the charger Hamiltonian rather than from intrinsically quantum effects [gyhm2022quantum]. Consequently, a fair comparison between different charging protocols requires fixing the operator norm of the charger Hamiltonian, thereby isolating the role of interaction-induced many-body dynamics.
Motivated by this perspective, we investigate how the interaction order $\kappa$, defined as the number of spins simultaneously coupled in the charger Hamiltonian, influences energy storage, instantaneous power, and quantum correlations under both unconstrained and norm-constrained (fair) charging conditions. Throughout this analysis, we consider a noninteracting battery described by Eq. (1), while the charger Hamiltonian [Eq. (2) with $J_{y}=J_{z}=0$] implements $\kappa$-body interactions.
Unconstrained Charging: Classical Scaling Effects
We first analyze the charging dynamics without imposing any constraint on the operator norm of the charger hamiltonian $\|\hat{H}_{C}\|$. This setting isolates classical contributions to energy storage arising purely from an increased energy scale of the charger, rather than from inherently quantum effects such as entanglement or multipartite correlations.
Figure 4: Fair charging with $\bm{h_{x}\neq 0}$ and $\bm{J_{x}=h_{x}}$. Instantaneous power $P_{i}$ as a function of time $t$ for a noninteracting battery with $h_{z}=1$, charged by a $\kappa$-local interacting charger in the presence of an additional transverse field of strength $h_{x}$. The charger Hamiltonian is normalized such that $\lVert\hat{H}_{C}\rVert=N(=8)$ by tuning the parameters $J_{x}$ and $h_{x}$ under the constraint $J_{x}=h_{x}$. The system size is $N=8$.
Figure 5: Fair charging with $\bm{h_{x}\neq 0}$ and $\bm{J_{x}>h_{x}}$. A noninteracting battery with $h_{z}=1$ is charged by a $\kappa=1,2,3$-local interacting charger in the presence of an additional transverse field of strength $h_{x}$. The charger Hamiltonian is normalized such that $\lVert\hat{H}_{C}\rVert=N(=8)$ by tuning the parameters under the constraint $J_{x}>h_{x}$. Panels (a) and (b) show, respectively, the instantaneous power $P_{i}$ and the ABEE $\overline{S}$ as functions of time $t$.
To show this classical advantage, we first consider the simplest case of parallel charging, where both the battery and the charger consist of noninteracting spins. The battery Hamiltonian is given by Eq. (1) with $h_{z}=1$, and the charger Hamiltonian is defined by Eq. (2) with $J_{x}=0$, supplemented by a transverse field $h_{x}$. The operator norm of the charger scales with the applied field: for $h_{x}=h_{z}$, we have $\lVert\hat{H}_{C}\rVert=N$, while for $h_{x}=N$, it increases as $\lVert\hat{H}_{C}\rVert=N^{2}$. As a consequence, the charging strength grows with $h_{x}$. Under $h_{x}=h_{z}$, the average stored energy reaches only half of the battery’s maximum capacity, $\Delta E_{\mathrm{max}}=Nh_{z}$, because the battery Hamiltonian partially counteracts the spin flips induced by the charger [shukla2025system]. Increasing $h_{x}$ further allows the charger to dominate the dynamics, coherently flipping all spins and achieving maximal energy storage, $\Delta E_{\mathrm{max}}=2Nh_{z}$ at $h_{x}=Nh_{z}$ [Fig. 3($a$)]. Importantly, this enhancement results entirely from the classical increase in energy scale, with no contribution from quantum correlations.
Next, we examine a fully interacting charger with $\kappa=N$ (all-to-all coupling) and $h_{x}=0$, while varying the interaction strength $J_{x}$. For $J_{x}=h_{z}$, the operator norm is small, $\lVert\hat{H}_{C}\rVert=1$, yielding modest maximum energy storage, $\Delta E_{\mathrm{max}}\approx 0.625$. By increasing the interaction to $J_{x}=Nh_{z}$, the operator norm rises to $\lVert\hat{H}_{C}\rVert=N$, resulting in $\Delta E_{\mathrm{max}}=Nh_{z}$, which matches the maximal stored energy achieved in the parallel scenario with $h_{x}=h_{z}$ [Fig. 3($b$)]. This comparison demonstrates that augmenting the charger strength either via the transverse field in parallel charging or via interaction strength in global charging produces equivalent increases in stored energy through classical effects alone.
The equivalence of these two protocols under unconstrained norms emphasizes that the observed enhancements in the unfair charging scenario are predominantly classical. Consequently, to assess genuine quantum advantages in charging performance, it is necessary to normalize the operator norm of the charger across different configurations. This fair charging condition ensures that any observed improvement in instantaneous power or energy transfer can be attributed to collective quantum effects, such as multipartite correlations or operator spreading, rather than trivial scaling of the Hamiltonian. Under this condition, one can systematically investigate the relationship between instantaneous power and quantum correlation measures such as quantum Fisher information and average bipartite entanglement entropy to reveal truly quantum contributions to the charging process.
Fair Charging with Flipping Term
To isolate genuine quantum effects in charging dynamics, we impose the fair charging condition by fixing the norm of the charger Hamiltonian, $\lVert\hat{H}_{C}\rVert=N$. This removes any classical advantage arising purely from differences in energy scale. Under this constraint, we study the role of the interaction order, $\kappa$, which defines the number of spins simultaneously coupled in the charger Hamiltonian, on the instantaneous power $P_{i}$.
We are fixing the charger Hamiltonina norm by considering $J_{x}=h_{x}$. For $\kappa=1$, corresponding to parallel charging of noninteracting spins, and increasing $\kappa$ to include multi-spin couplings ($\kappa=2,3,\ldots,N$), we find that $P_{i}$ initially grows, peaks, and then decreases. Importantly, with fair charging, increasing $\kappa$ alone does not increase the maximum instantaneous power beyond the parallel case [Fig. 4], indicating that multi-spin correlations by themselves are insufficient to provide a power advantage. This is consistent with the observation that classical enhancement dominates when the charger norm is unconstrained.
A different behavior emerges when the interaction strength $J_{x}$ becomes larger than the flipping term $h_{x}$. In this regime, $P_{i}$ reaches its peak more quickly, especially for larger $\kappa$ values [Fig. 5($a$)], demonstrating that cooperative multi-spin interactions can enhance charging power. To further understand this effect, we compare $P_{i}$ with the ABEE. Interestingly, BEE remains small at the time when $P_{i}$ peaks and increases only afterwards [Fig. 5($b$)], showing that the enhanced power arises not from entanglement growth but from cooperative many-body dynamics.
Under fair charging conditions, simply adding an external flipping term $h_{x}$ does not increase instantaneous power. Genuine enhancement occurs only when the interaction strength $J_{x}$ dominates, allowing cooperative dynamics among multiple spins to produce a measurable advantage. This highlights the critical role of strong interactions and system complexity in achieving quantum advantages in energy transfer.
Figure 6: Fair charging with $\bm{h_{x}=0}$. Noninteracting battery with $h_{z}=1$, charged by a $\kappa$-local interacting charger ($\kappa=1$ to NN) in the absence of a transverse field ($h_{x}=0$). The charger Hamiltonian is normalized as $\lVert\hat{H}_{C}\rVert=N(=8)$ by tuning the interaction strength $J_{x}$. (a) Instantaneous power $P_{i}$, (b) QFI, and (c) ABEE $\overline{S}$ as functions of time $t$, for different participation numbers $\kappa$, as indicated in panel (c). In panel (b), the horizontal dashed lines correspond to the entanglement bounds $\kappa(x)$, with $\kappa(1),\kappa(2),\dots$ indicated in the legend.
Fair Charging without the Flipping Term
To isolate the role of spin-spin interactions in the charging process, we eliminate single-spin rotations by setting the external flipping field to zero, $h_{x}=0$, in the charging Hamiltonian. This allows us to focus exclusively on interaction-driven dynamics and assess whether collective many-body effects alone can enhance charging performance. The Hamiltonian is normalized such that $\lVert\hat{H}_{C}\rVert=N$ for all interaction ranges, ensuring fair charging conditions.
In this interaction-only regime, the charging dynamics differ qualitatively from cases where local fields are present. The instantaneous power $P_{i}$ increases systematically with the interaction order $\kappa$, indicating that extending the number of spins participating simultaneously in the interaction promotes cooperative energy transfer. The enhancement becomes most pronounced in the global interaction limit $\kappa=N$, where all spins are collectively coupled. In this case, the maximum instantaneous power $P_{i}^{\mathrm{max}}$ reaches nearly $N$ times the value obtained under parallel charging [Fig. 6(a)]. This demonstrates that highly efficient charging can be achieved solely through collective interactions, even in the absence of local spin-flip processes.
To gain insight into the underlying quantum correlations, we analyze the QFI and the ABEE for all considered $\kappa$, while keeping $\lVert\hat{H}_{C}\rVert=N$ fixed. Both quantities increase with the interaction range and exhibit similar qualitative behavior [Figs. 6(b) and 6(c)], reflecting the growing multipartite nature of the quantum state as interactions become more global. Importantly, the peaks of the QFI and ABEE occur well after the instantaneous power reaches its maximum. This temporal mismatch indicates that multipartite entanglement is not the primary driver of peak power generation. Instead, collective interactions enable rapid energy transfer first, with entanglement building up subsequently as a consequence of the many-body dynamics.
The QFI provides a measure of the number of particles effectively participating in entanglement. In Fig. 6(b), the horizontal lines $\kappa(x)$ indicate specific participation numbers, and each time the QFI curve (solid line) crosses a line $\kappa(x)$, it signifies that $x+1$ particles are entangled. As the interaction order $\kappa$ increases, more particles participate in entanglement. For $\kappa<N$, only subsets of spins contribute, and the number of entangled particles does not grow monotonically with $\kappa$. Consequently, the instantaneous power increases with interaction order but not proportionally. In the fully connected regime ($\kappa=N$), all spins participate, resulting in a significantly larger instantaneous power compared to the $\kappa<N$ case. The temporal evolution of instantaneous power $P_{i}$, QFI, and ABEE also becomes closely aligned, reflecting fully collective and cooperative energy transfer throughout the system.
Fair Charging with Increasing Participation Number
Figure 7: Fair charging with increasing participation number: Noninteracting battery with $h_{z}=1$, charged by a 22-local interacting charger in the absence of a transverse field ($h_{x}=0$). The charger Hamiltonian is normalized as $\lVert\hat{H}_{C}\rVert=N(=8)$ by tuning the interaction strengths $J_{x_{1}}$ and $J_{x_{2}}$, corresponding to NN and NNN couplings, respectively, with $J_{x_{1}}>J_{x_{2}}$.
We now examine how extending the interaction range (pariticipation number) in the charger affects the charging dynamics under fair charging conditions. This is achieved by incorporating next-nearest-neighbor (NNN) couplings in the NN interaction, which increase the number of spins participating in the interaction without introducing fully global connectivity. The interaction part of the charging Hamiltonian is modified as
$\hat{H}_{\rm C}=J_{x_{1}}\sum_{j=1}^{N-1}\hat{\sigma}_{j}^{x}\hat{\sigma}_{j+1}^{x}+J_{x_{2}}\sum_{j=1}^{N-2}\hat{\sigma}_{j}^{x}\hat{\sigma}_{j+2}^{x},$ (11)
where $J_{x_{1}}$ and $J_{x_{2}}$ denote the strengths of NN and NNN interactions, respectively. To ensure a meaningful comparison between different interaction structures, we impose the fair charging condition by tuning the interaction strength, $J_{x_{1}}>J_{x_{2}}$, such that the operator norm of the charging Hamiltonian satisfies $\lVert\hat{H}_{C}\rVert=N$. This normalization eliminates any classical advantage arising from increased energy scales and isolates the effect of interaction geometry.
We compare the instantaneous power of chargers with NN interactions alone to those including both NN and NNN couplings, maintaining a fixed operator norm, $\lVert H_{C}\rVert=N$. The results indicate that adding NNN interactions decreases the maxima of instantaneous power $P_{i}$ compared to the NN-only protocol [Fig. 7]. This demonstrates that increasing the participation number through short-range interaction extensions does not automatically translate into enhanced charging power.
The observed suppression of power can be attributed to the emergence of competing many-body correlations induced by the additional interaction pathways. Rather than promoting coherent collective charging, the NNN couplings redistribute the injected energy across multiple channels, weakening the synchronization required for sharp power peaks. These results underscore that charging performance is governed not merely by the number of interacting spins, but by the coherence and structure of the interaction network. In particular, partially extended interaction schemes introduce correlation-induced competition effects that moderate collective energy transfer, in stark contrast to the strong power enhancement found in fully connected charging protocols.
IV Conclusion
In this work, we have investigated the physical mechanisms underlying energy storage and power enhancement in many-body quantum batteries, with particular emphasis on the role of quantum correlation and interaction structure, and the number of particles involved in entanglement. By systematically comparing energetic quantities such as stored energy and instantaneous charging power with a hierarchy of information-theoretic measures, we have clarified how entanglement and many-body dynamics jointly shape the charging process.
A central finding of our study is the robust temporal separation between energy transfer and correlation buildup. Across all battery–charger configurations considered, the instantaneous power consistently reaches its maximum before bipartite, tripartite, and multipartite entanglement attain their peak values. This establishes that peak charging is governed primarily by coherent collective transport enabled by interactions, rather than by fully developed entanglement or scrambling. Quantum correlations, while essential for stabilizing and redistributing energy at later times, do not directly determine the timing of maximal power output. This resolves an important ambiguity in the literature regarding the role of entanglement as a driver versus a byproduct of efficient charging.
Extending our analysis to $\kappa$-local charging protocols allowed us to disentangle genuine collective quantum effects from trivial classical scaling. Under unconstrained conditions, enhanced charging can be fully attributed to increases in the energy scale of the charger Hamiltonian. Imposing a fixed operator norm, thereby enforcing fair charging reveals that increasing the interaction order or the number of participating particles does not automatically lead to higher charging power. Charging performance is determined by the number of particles that are actually mutually correlated and contribute to entanglement. Fully collective interactions offer a genuine advantage, as all particles participate coherently in the charging process. In contrast, partially extended interaction schemes do not consistently increase the number of effectively interacting particles, and therefore, cannot guarantee improved charging efficiency.
Extending this analysis, adding NNN couplings in NN coupling under fair charging conditions further illustrates the limitations of partial interaction extensions. Compared to NN interactions alone, the maxima of instantaneous power are suppressed, confirming that merely increasing the participation number does not enhance charging performance.
Acknowledgments
RKS gratefully acknowledges Anupam and Sibasish Ghosh for their insightful and stimulating discussions.